OpenStudy (anonymous):

Choose the correct product of (2x - 4)2. 4x2 - 16 4x2 - 16x + 16 4x2 + 16 4x2 + 16x + 16 Medal and fan!! I think its A... but I might be wrong can someone please help!!

OpenStudy (anonymous):

\[(a-b)^2 = a^2 + b^2 - 2ab\] \[(2x-4)^2 = (2x)^2 + (4)^2 - 2*2x*4\]

OpenStudy (anonymous):

So would it be C? @sangya21

OpenStudy (kiakat):

No actually, it would be B. because both of the middle terms are negative so they don't cancel each other out.

OpenStudy (anonymous):

KiaKat is correct

OpenStudy (anonymous):

ohhh ok thanks :D

OpenStudy (kiakat):

No problem!